OpenStudy (justuu):

Medal and fan for help.. Use the Distributive Property to multiply 8 and 45 in your head. What is the answer? 32 + 40 + 72 40 + 40 = 80 320 + 40 = 360 320 + 5 = 325

OpenStudy (anonymous):

c

OpenStudy (justuu):

thank you

OpenStudy (anonymous):

yep

OpenStudy (justuu):

Use the Distributive Property to multiply 9 and 37 in your head. What is the answer? 270 + 7 = 277 27 + 63 = 90 30 + 63 = 93 270 + 63 = 333

OpenStudy (anonymous):

d

OpenStudy (justuu):

thank you can you help one more question?

OpenStudy (anonymous):

ya

OpenStudy (justuu):

Use the Distributive Property to multiply 9 and 37 in your head. After 37 is in expanded form, multiply 9 by each addend. (9 × 30) + (9 × 7) (9 × 30) + 7 30 + (9 × 7) 9 × 30 × 7